# Efficient Generative Model Training via Embedded Representation Warmup

### v1

@@ 5.5 Hyperparameter Tuning
  • To evaluate performance quickly, we reduce the sampling steps from the usual $250$ to $50$ and generate only $10\mathrm{k}$ samples (instead of $50\mathrm{k}$) to compute a preliminary FID score.
@@ 6 Conclusion and Future Work
In this work, we introduced Embedded Representation Initialization (ERW), a novel method that significantly enhances the training efficiency and representation quality of diffusion models by embedding high-quality representations from pretrained self-supervised encoders into the model’s representation processing region. Through both theoretical analysis and empirical validation, we demonstrated that ERW accelerates training convergence by up to 40$\times$ and improves the fidelity of generated data, establishing new state-of-the-art performance based on REPA [54]. Our findings underscore the importance of make representation embedded in initialization in optimizing diffusion model training.
Appendix A Loss Function Decomposition Proof
This section presents a theoretical decomposition of the diffusion model’s loss function into two orthogonal learning mechanisms: representation learning through latent-to-representation (L2R) mapping and generation inversion through representation-to-generation (R2G) dynamics. The decomposition reveals fundamental statistical independence between representation quality and generation fidelity in diffusion processes.
@@ Appendix A Loss Function Decomposition Proof
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|F_{\theta}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})\|^{2}\right],$ (19)
where $F_{\theta}(\mathbf{z}_{t},t)$ parameterizes the estimated score function, and $\lambda(t)$ is a time-dependent weighting function.
Our analysis hinges on the existence of a sufficient representation $\mathbf{r}_{t}=R(\mathbf{z}_{t})$ that statistically decouples the latent space geometry from generation dynamics. Formally, we assume:
@@
The representation $\mathbf{r}_{t}=R(\mathbf{z}_{t})$ is a sufficient statistic for $\mathbf{z}_{t}$ with respect to the data distribution $p(\mathbf{z}_{t})$, implying that $\mathbf{r}_{t}$ captures all necessary information about $\mathbf{z}_{t}$ for generation. This sufficiency leads to the decomposition of the Fisher score:
@@
where the generation score $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})$ and the representation score $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})$ are orthogonal under the Fisher-Rao metric:
@@
This orthogonality reflects the statistical independence between the generation mechanism and the representation learning process, ensuring that the two components can be optimized independently.
Under this assumption, the ground truth score decomposes into generation and representation components:
@@
Parametrizing the model’s score function as $F_{\theta}(\mathbf{z}_{t},t)=\mathcal{G}_{\theta}(\mathbf{r}_{t})+\mathcal{R}_{\theta}(\mathbf{z}_{t})$, where $\mathcal{G}_{\theta}$ and $\mathcal{R}_{\theta}$ independently model generation and representation scores, we expand the loss function:
$\mathcal{L}_{\text{diff}}$ $=\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|(\mathcal{G}_{\theta}(\mathbf{r}_{t})+\mathcal{R}_{\theta}(\mathbf{z}_{t}))-(\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})+\nabla\log p(\mathbf{r}_{t}))\|^{2}\right]$ (23)
$=\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{R}_{\theta}(\mathbf{z}_{t})-\nabla\log p(\mathbf{r}_{t})\|^{2}\right]$
$\quad+\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{G}_{\theta}(\mathbf{r}_{t})-\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})\|^{2}\right]$
$\quad+2\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\langle\mathcal{R}_{\theta}(\mathbf{z}_{t})-\nabla\log p(\mathbf{r}_{t}),\mathcal{G}_{\theta}(\mathbf{r}_{t})-\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})\rangle\right].$ (24)
The cross-term vanishes due to the orthogonality condition in Assumption 1, as:
@@
This yields the decoupled learning objectives:
$\mathcal{L}_{\text{diff}}=\underbrace{\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{R}_{\theta}(\mathbf{z}_{t})-\nabla\log p(\mathbf{r}_{t})\|^{2}\right]}_{\mathcal{L}_{\text{L2R}}}+\underbrace{\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{G}_{\theta}(\mathbf{r}_{t})-\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})\|^{2}\right]}_{\mathcal{L}_{\text{R2G}}}.$ (26)
Through the above decomposition, we effectively split the optimization objective of the diffusion model into two independent parts: Representation Learning $\mathcal{R}_{\theta}(\mathbf{z}_{t})$ and Generation Inversion $\mathcal{G}_{\theta}(\mathbf{r}_{t})$.
@@ B.1 Evaluation details
CKNNA (Centered Kernel Nearest-Neighbor Alignment) is a relaxed version of the popular Centered Kernel Alignment (CKA; Kornblith et al. 25) that mitigates the strict definition of alignment. We generally follow the notations in the original paper for an explanation [19].
@@
so this term only considers $k$-nearest neighbors at each $i$. In this paper, we randomly sample 10,000 images in the validation set in ImageNet [7] and report CKNNA with $k=10$ based on observation in Huh et al. [19] that smaller $k$ shows better a better alignment.
@@ B.2 Description of pretrained visual encoders
  • MoCov3 [6] studies empirical study to train MoCo [13, 5] on vision transformer and how they can be scaled up.
  • CLIP [42] proposes a contrastive learning scheme on large image-text pairs.
  • DINOv2 [39] proposes a self-supervised learning method that combines pixel-level and patch-level discriminative objectives by leveraging advanced self-supervised techniques and a large pre-training dataset.
@@ Appendix C Hyperparameter and More Implementation Details
Further implementation details. We implement our model based on the original SiT implementation [36]. Throughout the experiments, we use the exact same structure as DiT [40] and SiT [36]. We use AdamW [24, 35] with constant learning rate of 1e-4, $(\beta_{1},\beta_{2})=(0.9,0.999)$ without weight decay. To speed up training, we use mixed-precision (fp16) with a gradient clipping. We also pre-compute compressed latent vectors from raw pixels via stable diffusion VAE [44] and use these latent vectors. Because of this, we do not apply any data augmentation, but we find this does not lead to a big difference, as similarly observed in EDM2 [22]. We also use stabilityai/sd-vae-ft-ema decoder for decoding latent vectors to images. For MLP used for a projection, we use three-layer MLP with SiLU activations [9]. We provide a detailed hyperparameter setup in   Table 6.
Pretrained encoders. For MoCov3-B and -L models, we use the checkpoint in the implementation of RCG [32]; [^1] https://github.com/LTH14/rcg for other checkpoints, we use their official checkpoints released in their official implementations. To adjust a different number of patches between the diffusion transformer and the pretrained encoder, we interpolate positional embeddings of pretrained encoders.
Sampler. For sampling, we use the Euler-Maruyama sampler with the SDE with a diffusion coefficient $w_{t}=\sigma_{t}$. We use the last step of the SDE sampler as 0.04, and it gives a significant improvement, similar to the original SiT paper [36].
Training Tricks. We explore the influence of various training techniques on ERW’s performance. Notably, we observe performance improvements when incorporating Rotary Positional Embeddings [50].
Table 7: Impact of Training Tricks in ERW. Using the SD-VAE [44], ERW achieves an FID of 55.6 at 50K training steps on ImageNet class-conditional generation. This table illustrates how each training trick incrementally improves the FID, demonstrating that advanced design techniques enhance the original DiT performance.
@@
Discussion of Quantifiable Metrics.
In addition to the cost breakdown above, the following metrics are instrumental for rigorous quantitative analysis:
  • Iteration Time ($t_{\mathrm{iter}}$): The average time per training iteration (in seconds), which directly impacts $C_{w}$.
  • Warmup Iteration Count ($N_{w}$): The precise number of iterations employed during the warmup phase.
  • Main Training Iteration Count ($N_{m}$): The total number of iterations in the main training phase.
  • Hyperparameter Search Steps ($b_{i}$): The number of binary search steps needed for each hyperparameter.
  • Evaluation Efficiency: The use of a reduced evaluation protocol (e.g., FID measurement on 10K images using 50 sampling steps) accelerates the hyperparameter search process.
  • Overall Speedup Factor: The ratio $C_{m}/C_{w}$ is approximately 10, underscoring that the main training phase is the dominant cost.
@@ Appendix D Evaluation Details
We strictly follow the setup and use the same reference batches of ADM [8] for evaluation, following their official implementation. [^2] https://github.com/openai/guided-diffusion/tree/main/evaluations We use 8$\times$×NVIDIA H800 80GB GPUs or for evaluation and enable tf32 precision for faster generation, and we find the performance difference is negligible to the original fp32 precision.
@@
  • FID [15] measures the feature distance between the distributions of real and generated images. It uses the Inception-v3 network [51] and computes distance based on an assumption that both feature distributions are multivariate gaussian distributions.
  • sFID [37] proposes to compute FID with intermediate spatial features of the Inception-v3 network to capture the generated images’ spatial distribution.
  • IS [46] also uses the Inception-v3 network but use logit for evaluation of the metric. Specifically, it measures a KL-divergence between the original label distribution and the distribution of logits after the softmax normalization.
  • Precision and recall [27] are based on their classic definitions: the fraction of realistic images and the fraction of training data manifold covered by generated data.
@@ Appendix E Baselines
  • ADM [8] improves U-Net-based architectures for diffusion models and proposes classifier-guided sampling to balance the quality and diversity tradeoff.
  • VDM++ [23] proposes a simple adaptive noise schedule for diffusion models to improve training efficiency.
  • Simple diffusion [18] proposes a diffusion model for high-resolution image generation by exploring various techniques to simplify a noise schedule and architectures.
  • CDM [17] introduces cascaded diffusion models: similar to progressiveGAN [21], it trains multiple diffusion models starting from the lowest resolution and applying one or more super-resolution diffusion models for generating high-fidelity images.
  • LDM [44] proposes latent diffusion models by modeling image distribution in a compressed latent space to improve the training efficiency without sacrificing the generation performance.
  • U-ViT [2] proposes a ViT-based latent diffusion model that incorporates U-Net-like long skip connections.
  • DiffiT [12] proposes a time-dependent multi-head self-attention mechanism for enhancing the efficiency of transformer-based image diffusion models.
  • MDTv2 [11] proposes an asymmetric encoder-decoder scheme for efficient training of a diffusion-based transformer. They also apply U-Net-like long-shortcuts in the encoder and dense input-shortcuts in the decoder.
  • MaskDiT [55] proposes an asymmetric encoder-decoder scheme for efficient training of diffusion transformers, where they train the model with an auxiliary mask reconstruction task similar to MAE [14].
  • SD-DiT [56] extends MaskdiT architecture but incorporates self-supervised discrimination objective using a momentum encoder.
  • DiT [40] proposes a pure transformer backbone for training diffusion models based on proposing AdaIN-zero modules.
  • SiT [36] extensively analyzes how DiT training can be efficient by moving from discrete diffusion to continuous flow-based modeling.

### v2

@@ 5.5 Hyperparameter tuning
  2. (b)
    To evaluate performance quickly, we reduce the sampling steps from the usual 250250250 to 505050 and generate only $10\mathrm{k}$ samples (instead of $50\mathrm{k}$) to compute a preliminary FID score.
@@ 6 Conclusion and Future Work
In this work, we introduced Embedded Representation Warmup (ERW), a novel two-phase training framework that significantly enhances the training efficiency of diffusion models. By dedicating an initial phase to align the model’s early layers with a pretrained encoder, ERW establishes a strong semantic foundation that accelerates the subsequent generative training. Our key innovation is the explicit separation of representation alignment and generation, which, when combined with a decaying alignment schedule, proves more effective than continuous, single-phase regularization. We demonstrated empirically that ERW leads to substantial speedups in training convergence—up to 11.5$\times$ compared to REPA and achieves FID=1.41 with 350 epochs. Our ablations confirmed that targeting the early layers is crucial and that the two-phase approach is a cost-effective strategy for high-fidelity generative modeling.
Limitations and Future Work.
Despite its effectiveness, our work has limitations. First, the theoretical decomposition of the score function, while motivationally useful, is an idealization; the practical implementation relies on heuristic choices for the two-phase schedule and alignment targets. Second, the performance gains are sensitive to hyperparameters like the warmup duration and projection head depth, which currently require a manual tuning process. Future work will explore dynamically adjusting the alignment schedule based on training progress, rather than using a fixed decay.
Appendix A Loss Function Decomposition Analysis
This section presents a conceptual decomposition of the diffusion model’s loss function. Important note: This analysis is an idealization and relies on strong assumptions about representation sufficiency and orthogonality. In a real neural network, the components are not perfectly orthogonal, and the decomposition serves as a conceptual framework to motivate our method rather than a rigorous mathematical proof.
@@ Appendix A Loss Function Decomposition Analysis
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|F_{\bm{\theta}}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})\|^{2}\right],$ (17)
where $F_{\bm{\theta}}(\mathbf{z}_{t},t)$ parameterizes the estimated score function, and $\lambda(t)$ is a time-dependent weighting function.
Our analysis hinges on the idealized existence of a sufficient representation $\mathbf{r}_{t}=R(\mathbf{z}_{t})$ that could statistically decouple the latent space geometry from generation dynamics. Formally, we assume:
@@
The representation $\mathbf{r}_{t}=R(\mathbf{z}_{t})$ is assumed to be a sufficient statistic for $\mathbf{z}_{t}$ with respect to the data distribution $p(\mathbf{z}_{t})$. This implies that $\mathbf{r}_{t}$ would capture all necessary information about $\mathbf{z}_{t}$ for generation. This strong assumption leads to the hypothetical decomposition of the Fisher score:
@@
where the generation score $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})$ and the representation score $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})$ would be orthogonal under the Fisher-Rao metric:
@@
This idealized orthogonality would reflect the statistical independence between the generation mechanism and the representation learning process, ensuring that the two components could be optimized independently. However, in a practical neural network, these components are inevitably coupled.
Under this idealized assumption, the ground truth score decomposes into generation and representation components:
@@
Parametrizing the model’s score function as $F_{\bm{\theta}}(\mathbf{z}_{t},t)=\mathcal{G}_{\bm{\theta}}(\mathbf{r}_{t})+\mathcal{R}_{\bm{\theta}}(\mathbf{z}_{t})$, where $\mathcal{G}_{\bm{\theta}}$ and $\mathcal{R}_{\bm{\theta}}$ are intended to independently model generation and representation scores, we expand the loss function:
$\mathcal{L}_{\text{diff}}$ $=\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|(\mathcal{G}_{\bm{\theta}}(\mathbf{r}_{t})+\mathcal{R}_{\bm{\theta}}(\mathbf{z}_{t}))-(\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})+\nabla\log p(\mathbf{r}_{t}))\|^{2}\right]$ (21)
$=\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{R}_{\bm{\theta}}(\mathbf{z}_{t})-\nabla\log p(\mathbf{r}_{t})\|^{2}\right]$
$\quad+\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{G}_{\bm{\theta}}(\mathbf{r}_{t})-\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})\|^{2}\right]$
$\quad+2\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\langle\mathcal{R}_{\bm{\theta}}(\mathbf{z}_{t})-\nabla\log p(\mathbf{r}_{t}),\mathcal{G}_{\bm{\theta}}(\mathbf{r}_{t})-\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})\rangle\right].$ (22)
The cross-term would vanish due to the idealized orthogonality condition in Assumption 1, as:
@@
This yields the following decoupled learning objectives, which serve as a conceptual guide for our method:
$\mathcal{L}_{\text{diff}}\approx\underbrace{\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{R}_{\bm{\theta}}(\mathbf{z}_{t})-\nabla\log p(\mathbf{r}_{t})\|^{2}\right]}_{\mathcal{L}_{\text{L2R}}}+\underbrace{\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{G}_{\bm{\theta}}(\mathbf{r}_{t})-\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})\|^{2}\right]}_{\mathcal{L}_{\text{R2G}}}.$ (24)
Through the above conceptual decomposition, we can approximately split the optimization objective of the diffusion model into two largely independent parts: Representation Learning $\mathcal{R}_{\bm{\theta}}(\mathbf{z}_{t})$ and Generation Inversion $\mathcal{G}_{\bm{\theta}}(\mathbf{r}_{t})$. This provides the intuition for our two-phase training approach, even though perfect orthogonality is not achievable in practice.
@@ B.1 Evaluation details
CKNNA (Centered Kernel Nearest-Neighbor Alignment) is a relaxed version of the popular Centered Kernel Alignment (CKA; Kornblith et al. 2019) that mitigates the strict definition of alignment. We generally follow the notations in the original paper for an explanation (Huh et al., 2024).
@@
so this term only considers $k$-nearest neighbors at each $i$. In this paper, we randomly sample 10,000 images in the validation set in ImageNet (Deng et al., 2009) and report CKNNA with $k=10$ based on observation in Huh et al. (2024) that smaller $k$ shows better a better alignment.
@@ B.2 Description of pretrained visual encoders
  • MoCov3 (Chen et al., 2021) studies empirical study to train MoCo (He et al., 2020; Chen et al., 2020b) on vision transformer and how they can be scaled up.
  • CLIP (Radford et al., 2021) proposes a contrastive learning scheme on large image-text pairs.
  • DINOv2 (Oquab et al., 2024) proposes a self-supervised learning method that combines pixel-level and patch-level discriminative objectives by leveraging advanced self-supervised techniques and a large pre-training dataset.
@@ Appendix C Hyperparameter and More Implementation Details
Further implementation details. We implement our model based on the original SiT implementation (Ma et al., 2024). Throughout the experiments, we use the exact same structure as DiT (Peebles & Xie, 2023) and SiT (Ma et al., 2024). We use AdamW (Kingma, 2015; Loshchilov, 2017) with constant learning rate of 1e-4, $(\beta_{1},\beta_{2})=(0.9,0.999)$ without weight decay. To speed up training, we use mixed-precision (fp16) with a gradient clipping. We also pre-compute compressed latent vectors from raw pixels via stable diffusion VAE (Rombach et al., 2022) and use these latent vectors. Because of this, we do not apply any data augmentation, but we find this does not lead to a big difference, as similarly observed in EDM2 (Karras et al., 2024). We also use stabilityai/sd-vae-ft-ema decoder for decoding latent vectors to images. For MLP used for a projection, we use three-layer MLP with SiLU activations (Elfwing et al., 2018). We provide a detailed hyperparameter setup in   Table 6.
Pretrained encoders. For MoCov3-B and -L models, we use the checkpoint in the implementation of RCG (Li et al., 2024); [^1] https://github.com/LTH14/rcg for other checkpoints, we use their official checkpoints released in their official implementations. To adjust a different number of patches between the diffusion transformer and the pretrained encoder, we interpolate positional embeddings of pretrained encoders.
Sampler. For sampling, we use the Euler-Maruyama sampler with the SDE with a diffusion coefficient $w_{t}=\sigma_{t}$. We use the last step of the SDE sampler as 0.04, and it gives a significant improvement, similar to the original SiT paper (Ma et al., 2024).
Training Tricks. We explore the influence of various training techniques on ERW’s performance. Notably, we observe performance improvements when incorporating Rotary Positional Embeddings (Su et al., 2024).
Table 7: Impact of Training Tricks in ERW. Using the SD-VAE Rombach et al. (2022), ERW achieves an FID of 55.6 at 50K training steps on ImageNet class-conditional generation. This table illustrates how each training trick incrementally improves the FID, demonstrating that advanced design techniques enhance the original DiT performance.
@@ Appendix D Evaluation Details
We strictly follow the setup and use the same reference batches of ADM (Dhariwal & Nichol, 2021) for evaluation, following their official implementation. [^2] https://github.com/openai/guided-diffusion/tree/main/evaluations We use 8$\times$×NVIDIA H800 80GB GPUs or for evaluation and enable tf32 precision for faster generation, and we find the performance difference is negligible to the original fp32 precision.
@@
  • FID (Heusel et al., 2017) measures the feature distance between the distributions of real and generated images. It uses the Inception-v3 network (Szegedy et al., 2016) and computes distance based on an assumption that both feature distributions are multivariate gaussian distributions.
  • sFID (Nash et al., 2021) proposes to compute FID with intermediate spatial features of the Inception-v3 network to capture the generated images’ spatial distribution.
  • IS (Salimans et al., 2016) also uses the Inception-v3 network but use logit for evaluation of the metric. Specifically, it measures a KL-divergence between the original label distribution and the distribution of logits after the softmax normalization.
  • Precision and recall (Kynkäänniemi et al., 2019) are based on their classic definitions: the fraction of realistic images and the fraction of training data manifold covered by generated data.
@@ Appendix E Baselines
  • ADM (Dhariwal & Nichol, 2021) improves U-Net-based architectures for diffusion models and proposes classifier-guided sampling to balance the quality and diversity tradeoff.
  • VDM++ (Kingma & Gao, 2024) proposes a simple adaptive noise schedule for diffusion models to improve training efficiency.
  • Simple diffusion (Hoogeboom et al., 2023) proposes a diffusion model for high-resolution image generation by exploring various techniques to simplify a noise schedule and architectures.
  • CDM (Ho et al., 2022) introduces cascaded diffusion models: similar to progressiveGAN (Karras et al., 2018), it trains multiple diffusion models starting from the lowest resolution and applying one or more super-resolution diffusion models for generating high-fidelity images.
  • LDM (Rombach et al., 2022) proposes latent diffusion models by modeling image distribution in a compressed latent space to improve the training efficiency without sacrificing the generation performance.
  • U-ViT (Bao et al., 2023) proposes a ViT-based latent diffusion model that incorporates U-Net-like long skip connections.
  • DiffiT (Hatamizadeh et al., 2024) proposes a time-dependent multi-head self-attention mechanism for enhancing the efficiency of transformer-based image diffusion models.
  • MDTv2 (Gao et al., 2023) proposes an asymmetric encoder-decoder scheme for efficient training of a diffusion-based transformer. They also apply U-Net-like long-shortcuts in the encoder and dense input-shortcuts in the decoder.
  • MaskDiT (Zheng et al., 2024) proposes an asymmetric encoder-decoder scheme for efficient training of diffusion transformers, where they train the model with an auxiliary mask reconstruction task similar to MAE (He et al., 2022).
  • SD-DiT (Zhu et al., 2024) extends MaskdiT architecture but incorporates self-supervised discrimination objective using a momentum encoder.
  • DiT (Peebles & Xie, 2023) proposes a pure transformer backbone for training diffusion models based on proposing AdaIN-zero modules.
  • SiT (Ma et al., 2024) extensively analyzes how DiT training can be efficient by moving from discrete diffusion to continuous flow-based modeling.
  • REPA (Yu et al., 2024) proposes a representation alignment method for diffusion models by aligning the representation of the diffusion model with a pretrained encoder.